### v1

# positivity.bib

Abstract.
We investigate the positivity of self-adjoint polynomial Gaussian integral operators $\widehat{\kappa}_{\operatorname{PG}}$, that is, the multivariable kernel $\kappa_{\operatorname{PG}}$ is a product of a polynomial $P$ and a Gaussian kernel $\kappa_{G}$.
We show that $\widehat{\kappa}_{\operatorname{PG}}$ can be only positive if the Gaussian part is positive. This has an important corollary for the bipartite entanglement of the density operators $\widehat{\kappa}_{\operatorname{PG}}$: if the Gaussian density operator $\widehat{\kappa}_{G}$ fails the Peres–Horodecki criterion, then the corresponding polynomial Gaussian density operators $\widehat{\kappa}_{\operatorname{PG}}$ also fail the criterion for all $P$, hence they are all entangled.
We introduce a new preorder $\preceq$ on Gaussian kernels such that if $\kappa_{G_{0}}\preceq\kappa_{G_{1}}$ then $\widehat{\kappa}_{\operatorname{PG}_{0}}\geq 0$ implies $\widehat{\kappa}_{\operatorname{PG}_{1}}\geq 0$ for all polynomials $P$. Therefore, deciding the positivity of a polynomial Gaussian operator determines the positivity of a lot of another polynomial Gaussian operators which have the same polynomial factor. This preorder has implication for the entanglement problem, too.
We also prove that polynomial Gaussian operators with polynomials of odd degree cannot be positive semidefinite.
Key words and phrases:
entanglement, separability, positive semidefinite operators, integral operators, Hilbert-Schmidt operators, Gaussian operators, polynomial Gaussian, density operators, quantum theory, phase space, position representation, symplectic transformation
The first author was supported by the National Research, Development and Innovation Office – NKFIH, grants no. 124749 and 146922, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
The second author was supported by the Hungarian National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary grants no. 2022-2.1.1-NL-2022-00004 and 134437.
The third author thanks the ”Frontline” Research Excellence Programme of the NKFIH (Grant no. KKP133827) and Project no. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
1. Introduction
1.1. History and motivation
The concept of self-adjoint integral operators has been in the center of mathematical and physical research over the last century. Fredholm, Hilbert, Mercer, and Schmidt were the first ones who established the cornerstones of this area [Fredholm1903, Hilbert1904, Mercer, schmidt1905entwickelung, Schmidt1907, Stewart], which is still an active field of research, see the monographs [Arvind1995, Folland+1989, Gosson_Harmonic, deGosson_book, Gosson_trace_class, Gosson]. The foundation of quantum mechanics in the $1920$s brought even more attention to the notions of integral operators and infinite dimensional Hilbert spaces. One of the most important mathematical objects of quantum information theory and quantum physics is the density operator which is a positive self-adjoint trace class operator with trace one [Neumann].
Technically, density operators acting on infinite dimensional Hilbert spaces can be represented by $L^{2}$ kernels in position representation, or equivalently by Wigner functions in phase space, or with characteristic functions in another phase space [Serafini, nielsen_chuang_2010]. Phase space methods were first applied by Weyl, Wigner, Husimi, and Moyal [Weyl, Wigner, Husimi, Moyal], and they found many applications in mathematics, quantum chemistry, statistical mechanics, quantum optics and quantum information theory [Folland+1989, Hillery, Lee, Schleich, Weinbub, Salazar, Adesso].
Despite the large literature on integral operators, it is still a very hard problem to determine the spectrum, and even the positivity of operators in concrete cases. The positivity check of different models of quantum mechanics has already started in the $1960$s with the so-called KLM conditions [Kastler, Loupias1, Loupias2], and later further studies of trace class operators were carried out [Gosson_trace_class, Narkowich1, Narkowich2, Werner, Luef]. It is still an active area of research to check the positivity of an operator without actually calculating the full spectrum [Gosson, Newton]. The dynamics of the density operator of a subsystem are usually given by master equations, where evolution in time is governed by partial differential equations. Usually, a quantum-mechanical system interacts with external quantum systems and these interactions significantly change the dynamics of the examined subsystem, causing quantum dissipation and decoherence [book1]. A good example for this phenomenon is the quantum Brownian motion [HPZ, H-Y], where the master equations are derived from first principles and they may lead to positivity violations of the density operator at a later time depending on the external parameters of the model [Gnutzmann, BLH, HCSB, HBCSCS].
At the beginning of modern quantum mechanics a new phenomenon, the quantum entanglement and its complementary notion, the separability emerged [EPR, Bell], and they still pose serious challenges to both mathematicians and theoretical physicists [Open, Zyczkowski]. In two breakthrough papers, Peres [Perescrit] and the Horodecki family [HORODECKI19961] independently found a necessary condition for bipartite separability by checking positivity after the so-called partial transpose operation, which provides a very important link between entanglement and positivity. Another important result is due to Werner and Wolf, who proved (based on a previous work of Simon) that the Peres–Horodecki criterion is a necessary and sufficient condition of separability for all bipartite $1$ versus $n$ modes Gaussian quantum states [WernerWolf, SimonR, Duan, Lami]. Determining entanglement of non-Gaussian density operators in infinite dimensional Hilbert spaces is an extremely difficult task [Pirandola, Vogel, Miki, Volume]. Another approach to the entanglement problem is using so-called witness operators given by Gühne and Tóth [GUHNE]. The situation becomes even more complicated if we consider the time evolution of the entanglement of composite quantum systems [nielsen_chuang_2010, kina_entanglement].
Gaussian kernels occur quite frequently in physics, for example coupled harmonic oscillators in thermal distribution can be described by Gaussian operators of the form $\widehat{\rho}=\exp{\big{(}-\beta\widehat{H}_{\text{osc}}\big{)}}/\operatorname{Tr}\big{[}\exp{\big{(}-\beta\widehat{H}_{\text{osc}}\big{)}}\big{]}$, if the Hamiltonian operator $\widehat{H}_{\text{osc}}$ is at most quadratic in position and momentum operators [Plenio_2004]. The positivity of a Gaussian kernel is fully determined by its covariance matrix (see [Serafini]), while the behaviour of non-Gaussian kernels is more subtle. Our main goal is to consider more general polynomial Gaussian integral operators, that is, operators whose kernels can be written as a polynomial multiple of a Gaussian, see (1.8) for the precise definition. Contrary to the Gaussian case, for polynomial Gaussian operators no finite positivity test is known. Polynomial Gaussian forms appear naturally in physics. Examples in quantum optics can be found in [meng2023entangled]. Excited states of coupled oscillators in position representation have also this form: polynomial Gaussian operators can be obtained from those excited states by mixing some of them with positive weights [kina_entanglement, Dattoli1996, PhysRevA.54.5378]. Studying the entanglement of the above composite system with the Peres–Horodecki criterion translates to checking the positivity of a polynomial Gaussian operator. Our strategy is to prove results about the positivity of polynomial Gaussian operators, which can be translated to the language of entanglement by the Peres–Horodecki criterion. The entanglement problems have already found several applications in various disciplines e. g. quantum chemistry [Ding2021], quantum information theory [Coladangelo2020, Infinite], quantum optics [meng2023entangled], quantum communication [Piveteau2022] and quantum computers [Fauseweh2024]. A recent mathematical breakthrough [MIPRE, MIPRE2] in the area of quantum computational complexity provides connection to the areas of quantum information, operator algebras, and approximate representation theory; in particular it solves Tsirelson’s problem [Tsirelson_Bell] from quantum information theory, and the equivalent Connes’ embedding problem [Connes] from the theory of von Neumann algebras. For the above problems see also [Junge], where it is stated that “finite dimensional quantum models do not suffice to describe all bipartite correlations”, emphasizing the role of operators acting on infinite dimensional Hilbert spaces. This wide applicability and connections between different areas served as a source of inspiration for us, too.
1.2. Theoretical background
We denote by $L^{2}\left(\mathbb{R}^{n}\right)$ the Hilbert space of the complex-valued square integrable functions defined on $\mathbb{R}^{n}$ with the scalar product [^1] In physics the usual convention is $\langle f,g\rangle=\int_{\mathbb{R}^{n}}f^{*}(x)g(x)\,\mathrm{d}x$. $\langle f,g\rangle=\int_{\mathbb{R}^{n}}f(x)g^{*}(x)\,\mathrm{d}x$, where $z^{*}$ denotes the complex conjugate of $z$. A kernel $\kappa\in L^{2}(\mathbb{R}^{2n})$ defines a Hilbert-Schmidt integral operator $\widehat{\kappa}\colon L^{2}\left(\mathbb{R}^{n}\right)\to L^{2}\left(\mathbb{R}^{n}\right)$ by the formula
(1.1) $\left(\widehat{\kappa}f\right)(x)=\int_{\mathbb{R}^{n}}\,\kappa(x,y)f(y)\,\mathrm{d}y,$
see e. g. [Simon, Stein]. We call $\widehat{\kappa}$ self-adjoint if $\widehat{\kappa}=\widehat{\kappa}^{\dagger}$. For continuous kernels $\kappa$ this is equivalent to the property that $\kappa(y,x)=\kappa^{*}(x,y)$ for all $x,y\in\mathbb{R}^{n}$. We say that $\widehat{\kappa}$ is positive semidefinite if $\langle\widehat{\kappa}f,f\rangle\geq 0$ for all $f\in L^{2}(\mathbb{R}^{n})$; we also use the notation $\widehat{\kappa}\geq 0$. If $\widehat{\kappa}$ is positive semidefinite, it is necessarily self-adjoint. Every Hilbert-Schmidt integral operator is compact, that is, the closure of the image of the open unit ball under the operator is compact, see e. g. [Rudin]. Therefore, it has only countably many eigenvalues $\{\lambda_{i}\}^{\infty}_{i=0}$, see [Rudin, Theorem 4.25]. The eigenvalue equation of $\widehat{\kappa}$ is a Fredholm-type integral equation
(1.2) $\int_{\mathbb{R}^{n}}\,\kappa(x,y)\phi_{i}(y)\,\mathrm{d}y=\lambda_{i}\phi_{i}(x),\quad\text{where }\phi_{i}\in L^{2}(\mathbb{R}^{n}).$
We say that the operator $\widehat{\kappa}$ is trace class if $\sum_{i=0}^{\infty}|\lambda_{i}|<\infty$; in this case we define the trace of $\widehat{\kappa}$ as
(1.3) $\operatorname{Tr}(\widehat{\kappa})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{\infty}\lambda_{i}.$
If $\kappa$ is continuous, then (see [Brislawn] and the original [Duflo]) we have the formula
(1.4) $\operatorname{Tr}(\widehat{\kappa})=\int_{\mathbb{R}^{n}}\kappa(x,x)\,\mathrm{d}x.$
The Schwartz space $\mathcal{S}(\mathbb{R}^{n})$ is the set of rapidly decreasing smooth functions defined as follows. A smooth function $f\colon\mathbb{R}^{n}\to\mathbb{C}$ satisfies $f\in\mathcal{S}(\mathbb{R}^{n})$ if for all multiindices $\alpha=(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}$ and $\beta=(\beta_{1},\dots,\beta_{n})\in\mathbb{N}^{n}$ we have
$\sup_{x\in\mathbb{R}^{n}}|x^{\alpha}(D^{\beta}f)(x)|<\infty,$
where we use the notation $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ and $D^{\beta}=\partial_{1}^{\beta_{1}}\cdots\partial_{n}^{\beta_{n}}$; for more on Schwartz functions see e. g. [SR, Section V.3]. If $\kappa\in\mathcal{S}(\mathbb{R}^{2n})$ is a Schwartz kernel, then $\widehat{\kappa}$ is a trace class operator, see [Gosson_Harmonic, Proposition 287] or [Brislawn, Proposition 1.1] with the remark afterwards. Note that in quantum mechanical descriptions of physical systems our eigenvalues are probabilities. Thus, we say that $\widehat{\rho}$ is a density operator if it is positive semidefinite with $\operatorname{Tr}(\widehat{\rho})=1$, so its eigenvalues satisfy $\lambda_{i}\geq 0$ and $\sum_{i=0}^{\infty}\lambda_{i}=1$. For more on Hilbert-Schmidt, compact, and trace class operators see e. g. [Gosson_Harmonic, deGosson_book, Stein]. Let $\kappa\in\mathcal{S}(\mathbb{R}^{2n})$ be a continuous kernel such that $\widehat{\kappa}$ is trace class. The following theorem basically dates all the way back to Mercer [Mercer].
Theorem 1.1 (Mercer).
Assume that $\kappa\in L^{2}(\mathbb{R}^{2n})$ is a continuous kernel. Then $\widehat{\kappa}$ is positive semidefinite if and only if for all $x_{1},\dots,x_{k}\in\mathbb{R}^{n}$ and $c_{1},\dots,c_{k}\in\mathbb{C}$ we have
(1.5) $\sum_{i,j=1}^{k}c_{i}c^{*}_{j}\kappa(x_{i},x_{j})\geq 0.$
For an operator $\widehat{\kappa}$ the Wigner–Weyl transform connects the position representation $\kappa(x,y)$ and the phase space representation [^2] If $\widehat{\kappa}$ is a density operator, then $W(x,p)$ is called Wigner function. $W(x,p)$ such that for all $x,y,p\in\mathbb{R}^{n}$ we have
(1.6) $W(x,p)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}\exp{\left\{-ip^{\text{T}}y\right\}}\kappa\left(x+\frac{y}{2},x-\frac{y}{2}\right)\,\mathrm{d}y,$
and the inverse transform is
(1.7) $\kappa(x,y)=\int_{\mathbb{R}^{n}}W\left(\frac{x+y}{2},p\right)\exp{\left\{ip^{\text{T}}\left(x-y\right)\right\}}\,\mathrm{d}p;$
for more on these concepts the reader might consult the monograph [Folland+1989]. Note that here we used the convention that the Planck-constant satisfies $\hbar=1$.
1.3. Statement of results
Let $A,B,C\in\mathbb{R}^{n\times n}$ such that $A$ and $C$ are positive definite and $B$ is arbitrary. We consider the Gaussian kernel $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ as
$\kappa_{G}(x,y)=\exp\left\{-(x-y)^{\text{T}}A(x-y)-i(x-y)^{\text{T}}B(x+y)-(x+y)^{\text{T}}C(x+y)\right\}.$
Let $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a polynomial in $2n$ variables and define the polynomial Gaussian kernel [^3] The same terminology was used in [Roux_polygaussian] in a similar context. $\kappa_{\operatorname{PG}}$ as [^4] We could have added the linear term $iD^{\text{T}}(x-y)+E^{\text{T}}(x+y)$ to the exponent of $\kappa_{G}$ with some given vectors $D,E\in\mathbb{R}^{n}$. However, $D$ does not change the spectrum of $\widehat{\kappa}_{\operatorname{PG}}$, and $E$ does not change its positivity, so they are irrelevant to us.
(1.8) $\kappa_{\operatorname{PG}}(x,y)=P(x,y)\kappa_{G}(x,y).$
Since $\kappa_{\operatorname{PG}}\in\mathcal{S}(\mathbb{R}^{2n})$, the operator $\widehat{\kappa}_{\operatorname{PG}}$ is trace class. Note that if $\widehat{\kappa}_{\operatorname{PG}}\geq 0$ then $P$ must be self-adjoint, that is, $P(y,x)=P^{*}(x,y)$ for all $x,y\in\mathbb{R}^{n}$.
1.3.1. When the Gaussian is not positive
The following theorem is the main result of our paper, stating that if a Gaussian operator is not positive semidefinite, then the corresponding polynomial Gaussians cannot be positive semidefinite, either.
Theorem 1.2.
Let $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a non-zero polynomial and let $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a Gaussian kernel. If $\widehat{\kappa}_{G}$ is not positive semidefinite, then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite, either.
1.3.2. When the polynomial is of odd degree
We show that if a polynomial $P$ is of odd degree, then no corresponding polynomial Gaussian operator can be positive semidefinite.
Theorem 1.3.
Assume that $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ is a polynomial of odd degree, and $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ is any Gaussian kernel. Then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite.
The above theorem generalizes [Newton, Proposition 3.1], where the significantly easier $n=\deg P=1$ case was settled. It also explains why polynomial Gaussian operators with polynomials of odd degree do not really appear in physics.
Let us say that $P$ is reducible to odd degree if there is an index set $I\subset\{1,\dots,n\}$ such that substituting $x_{i}=y_{i}=0$ for all $i\in I$ into $P$ transforms $P$ into a polynomial of odd degree in $2(n-|I|)$ variables. Theorem 1.1 easily implies that applying this transformation to a kernel maps a positive semidefinite operator into a positive semidefinite one. Therefore, Theorem 1.3 immediately yields the following generalization.
Theorem 1.4.
Let $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a polynomial which is reducible to odd degree, and let $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a Gaussian kernel. Then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite.
1.3.3. A useful preorder on Gaussian kernels
We introduce a relation which allows us to vary the Gaussian part in polynomial Gaussian (and more general) kernels. Fix a positive integer $n$, and consider matrices $A,B,C\in\mathbb{R}^{n\times n}$ such that both $A$ and $C$ are symmetric. Define self-adjoint Gaussian functions $\theta=\theta(A,B,C)\colon\mathbb{R}^{2n}\to\mathbb{C}$ as
$\theta(x,y)=\exp\left\{-(x-y)^{\text{T}}A(x-y)-i(x-y)^{\text{T}}B(x+y)-(x+y)^{\text{T}}C(x+y)\right\}.$
Clearly $\theta$ is a kernel in $L^{2}(\mathbb{R}^{2n})$ if and only if both $A$ and $C$ are positive definite. We want to address the following problem.
Problem 1.5.
Let $\sigma\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a given self-adjoint function and assume that $\sigma\theta(A_{i},B_{i},C_{i})\in L^{2}(\mathbb{R}^{2n})$ for $i\in\{0,1\}$. When is it true that
$\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0\quad\Longrightarrow\quad\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0?$
The following definition and theorem answer this problem completely.
Notation 1.6.
Let $I_{n}\in\mathbb{R}^{n\times n}$ denote the identity matrix. Let $\mathcal{G}(n)$ be the set of triples $(A,B,C)$ such that $A,B,C\in\mathbb{R}^{n\times n}$ and $A,C$ are symmetric. Let
$\mathcal{G}^{+}(n)=\left\{(A,B,C)\in\mathcal{G}(n):A,C\text{ are positive definite and }\widehat{\theta}(A,B,C)\geq 0\right\}.$
Definition 1.7.
We define a relation $(\mathcal{G},\preceq)$ on $\mathcal{G}(n)$ as follows. For two triples $(A_{0},B_{0},C_{0}),(A_{1},B_{1},C_{1})\in\mathcal{G}(n)$ we write $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$ if there exists $r\geq 0$ such that
$(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})\in\mathcal{G}^{+}(n).$
For two Gaussian kernels $\kappa_{0}$ and $\kappa_{1}$ we write $\kappa_{0}\preceq\kappa_{1}$ if their defining matrix triples satisfy $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$.
Remark 1.8.
In Definition 1.7 the matrices $A_{1}-A_{0}$ and $C_{1}-C_{0}$ are symmetric, so for large enough $r$ the matrices $A_{1}-A_{0}+rI_{n}$ and $C_{1}-C_{0}+rI_{n}$ are both positive definite, yielding ${\theta}(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})\in L^{2}(\mathbb{R}^{2n})$. Theorem 1.15 will imply that if $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$, then any choice of $r\geq 0$ witnesses it as long as the matrices $A_{1}-A_{0}+rI_{n}$ and $C_{1}-C_{0}+rI_{n}$ are positive definite. Hence checking the relation $\preceq$ requires to determine the positivity of a single Gaussian operator, which can be easily done by calculating the eigenvalues of a $2n\times 2n$ matrix as we will see in Claim 2.4.
The following theorem states that this is precisely the right definition for us.
Theorem 1.9.
For $(A_{0},B_{0},C_{0}),(A_{1},B_{1},C_{1})\in\mathcal{G}(n)$ the following are equivalent:
  1. (1)
    $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$;
  2. (2)
    for each self-adjoint $\sigma\colon\mathbb{R}^{2n}\to\mathbb{C}$ if $\sigma\theta(A_{i},B_{i},C_{i})\in L^{2}(\mathbb{R}^{2n})$ for $i\in\{0,1\}$ and $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0$, then $\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0$;
  3. (3)
    for every self-adjoint Gaussian kernel $\sigma$ if $\sigma\theta(A_{i},B_{i},C_{i})\in L^{2}(\mathbb{R}^{2n})$ for $i\in\{0,1\}$ and $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0$, then $\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0$.
The next corollary shows how to use this theorem in practice. It states that deciding the positivity of a single operator determines the positivity of a lot of other operators as well.
Corollary 1.10.
Let a self-adjoint function $\sigma\colon\mathbb{R}^{2n}\to\mathbb{C}$ and $(A_{0},B_{0},C_{0})\in\mathcal{G}(n)$ be given such that $\sigma\theta(A_{0},B_{0},C_{0})\in L^{2}(\mathbb{R}^{2n})$, and consider all triples $(A_{1},B_{1},C_{1})$ for which $\sigma\theta(A_{1},B_{1},C_{1})\in L^{2}(\mathbb{R}^{2n})$. Assume that we have already determined the positivity of $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})$, then there are two possibilities:
  1. (i)
    $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0\Rightarrow\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0\text{ for all }(A_{1},B_{1},C_{1})\succeq(A_{0},B_{0},C_{0})$;
  2. (ii)
    $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\not\geq 0\Rightarrow\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\not\geq 0\text{ for all }(A_{1},B_{1},C_{1})\preceq(A_{0},B_{0},C_{0})$.
Theorem 1.9 (and directly Fact 4.1) imply that the relation $\preceq$ is transitive, and clearly reflexive, which makes it a preorder. The following claim describes the polynomials which define positive operators with any positive Gaussian factor.
Claim 1.11.
We have $(I_{n},0,I_{n})\preceq(A,B,C)$ for all $(A,B,C)\in\mathcal{G}^{+}(n)$. Hence for a self-adjoint polynomial $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ the following are equivalent:
  1. (1)
    $\widehat{P\theta}(A,B,C)\geq 0$ for all $(A,B,C)\in\mathcal{G}^{+}(n)$;
  2. (2)
    $\widehat{P\theta}(I_{n},0,I_{n})\geq 0$;
  3. (3)
    $\sum_{i,j=1}^{k}c_{i}c^{*}_{j}P(x_{i},x_{j})\geq 0$ for all $x_{1},\dots,x_{k}\in\mathbb{R}^{n}$ and $c_{1},\dots,c_{k}\in\mathbb{C}$.
As it is much easier to check (3) than (1), the above equivalence is especially useful. We demonstrate this by the following example.
Example 1.12.
Let $P\colon\mathbb{R}^{2}\to\mathbb{R}$ defined as $P(x,y)=x^{\ell}y^{m}+x^{m}y^{\ell}$ with $\ell,m\in\mathbb{N}$. Then $P$ satisfies Claim (1.11) (1) if and only if $\ell=m$. Indeed, if $\ell=m$ then (3) is clearly satisfied. Assume to the contrary that $\ell>m$. Applying (3) with $c_{1}=2,c_{2}=-1$, $x_{1}=x$, $x_{2}=y$ and using the notation $\lambda=\frac{y}{x}$ we obtain that
(1.9) $0\leq 8x^{\ell+m}-4(x^{\ell}y^{m}+x^{m}y^{\ell})+2y^{\ell+m}=2x^{\ell+m}(\lambda^{\ell}-2)(\lambda^{m}-2).$
We can choose $x,y>0$ such that $\lambda^{m}<2<\lambda^{\ell}$, which contradicts (1.9).
Corollary 1.13.
Let $(A_{i},B_{i},C_{i})\in\mathcal{G}(n)$, $i\in\{0,1\}$. Then
  1. (1)
    $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$Rightarrow $A_{1}-C_{1}\geq A_{0}-C_{0}$,
  2. (2)
    $A_{1}-C_{1}\geq A_{0}-C_{0}$ and $B_{1}-B_{0}$ is symmetric $\Rightarrow$ $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$.
Notation 1.14.
Let $(A_{i},B_{i},C_{i})\in\mathcal{G}(n)$, $(i=0,1)$. We define the equivalence relation $(A_{0},B_{0},C_{0})\approx(A_{1},B_{1},C_{1})$ such that
$(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})\text{ and }(A_{1},B_{1},C_{1})\preceq(A_{0},B_{0},C_{0}).$
The next theorem characterizes the equivalence of the triples $(A,B,C)\in\mathcal{G}(n)$.
Theorem 1.15.
Let $(A_{i},B_{i},C_{i})\in\mathcal{G}(n)$, $i\in\{0,1\}$. The following are equivalent:
  1. (1)
    $(A_{0},B_{0},C_{0})\approx(A_{1},B_{1},C_{1})$,
  2. (2)
    $A_{1}-C_{1}=A_{0}-C_{0}$ and $B_{1}-B_{0}$ is symmetric.
1.3.4. Entanglement
Here we enclose an introduction of entanglement, then we deduct some consequences of our earlier positivity results. A physical system in quantum mechanics can be described by a density operator $\widehat{\rho}\colon\mathcal{H}\rightarrow\mathcal{H}$ acting on a complex Hilbert space $\mathcal{H}$, see e. g. John von Neumann’s original works [Neumann, Neumann1927]. Here we will consider the Hilbert space of square integrable complex functions $\mathcal{H}=L^{2}\left(\mathbb{R}^{n}\right)$, where $n$ is the number of degrees of freedom in our physical system. Our indices $1,\dots,n$ correspond to variables which might, for example, represent the location or moment of some particles. We can divide our system to $m$ physical subsystems by partitioning our index set $\{1,\dots,n\}$ into $m$ pairwise disjoint sets $P_{1},\dots,P_{m}$ of sizes $d_{1},\dots,d_{m}$, respectively, where $\sum^{m}_{k=1}d_{k}=n$. For any $k\in\{1,\dots,m\}$ let $\mathcal{H}_{k}=L^{2}\left(\mathbb{R}^{P_{k}}\right)\cong L^{2}(\mathbb{R}^{d_{k}})$, which is the Hilbert space belonging to the $k$th physical subsystem. This partition decomposes $\mathcal{H}$ to the tensor product
(1.10) $\mathcal{H}=\bigotimes_{k=1}^{m}\mathcal{H}_{k}.$
For more information on $C^{*}$-algebras see e. g. [Murphy].
Definition 1.16.
Let $\widehat{\rho}\colon L^{2}\left(\mathbb{R}^{n}\right)\to L^{2}\left(\mathbb{R}^{n}\right)$ be a density operator and assume that $m$ physical subsystems are given by an $m$-element partition $\mathcal{P}=\{P_{1},\dots,P_{m}\}$ of $\{1,\dots,n\}$, and let $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ be the corresponding Hilbert spaces. We call $\widehat{\rho}$ separable with respect to $\mathcal{P}$ if it can be written as
(1.11) $\widehat{\rho}=\sum_{j=1}^{\infty}\pi_{j}\bigotimes_{k=1}^{m}\widehat{\rho}^{\,(k)}_{j},\text{ where }\pi_{j}\geq 0\text{ and }\sum_{j=1}^{\infty}\pi_{j}=1,$
and $\widehat{\rho}^{\,(k)}_{j}\colon\mathcal{H}_{k}\to\mathcal{H}_{k}$ are density operators for all $j\geq 1$. We say that $\widehat{\rho}$ is entangled with respect to $\mathcal{P}$ if it cannot be written as (1.11).
Definition 1.17.
Let $\mathcal{P}$ and $\mathcal{P}^{\prime}$ be two partitions of $\{1,\dots,n\}$. We say that $\mathcal{P}$ is a refinement of $\mathcal{P}^{\prime}$ if for all $P\in\mathcal{P}$ there is a $P^{\prime}\in\mathcal{P}^{\prime}$ such that $P\subset P^{\prime}$.
The definition of separability easily implies the following, which allows us to prove multipartite entanglement by proving bipartite entanglement for a coarser two-element partition.
Fact 1.18.
Assume that partition $\mathcal{P}$ is a refinement of $\mathcal{P}^{\prime}$. If the density operator $\widehat{\rho}$ is separable with respect to $\mathcal{P}$, then it is separable with respect to $\mathcal{P}^{\prime}$. Equivalently, if $\widehat{\rho}$ is entangled with respect to $\mathcal{P}^{\prime}$, then it is entangled with respect to $\mathcal{P}$, too.
From now on, we only consider the bipartite case, that is, a $2$-element partition $\mathcal{P}=\{P_{1},P_{2}\}$ of our index set $\{1,\dots,n\}$ with given sizes $d_{1},d_{2}$. We consider the separability of density operators $\widehat{\rho}$ with respect to $\mathcal{P}$. By varying the indices if necessary, we may assume without loss of generality that $P_{1}=\{1,\dots,d_{1}\}$ and $P_{2}=\{d_{1}+1,\dots,n\}$. Recall that the Wigner–Weyl transform (1.6) and its inverse (1.7) provide the connection between the phase space and the position representations. If $\widehat{\rho}$ is separable with respect to $\mathcal{P}$, similarly to (1.11) it can be written in phase space representation as
$W(x,p)=W\big{(}x^{(1)},p^{(1)};x^{(2)},p^{(2)}\big{)}=\sum_{j=1}^{\infty}\pi_{j}W^{(1)}_{j}\big{(}x^{(1)},p^{(1)}\big{)}\cdot W^{(2)}_{j}\big{(}x^{(2)},p^{(2)}\big{)},$
where $\pi_{j}\geq 0$ and $\sum_{j=1}^{\infty}\pi_{j}=1$, and $W^{(i)}_{j}$ is the phase space representation of the density operator $\widehat{\rho}^{\,(i)}_{j}$ coming from Definition 1.16 for all $i\in\{1,2\}$ and $j\geq 1$, and we used the notation
$x$ $=(x_{1},\dots,x_{n})^{\text{T}},\leavevmode\nobreak\ x^{(1)}=\left(x_{1},\dots,x_{d_{1}}\right)^{\text{T}},\text{ and }x^{(2)}=\left(x_{d_{1}+1},\dots,x_{n}\right)^{\text{T}};$
$p$ $=(p_{1},\dots,p_{n})^{\text{T}},\leavevmode\nobreak\ p^{(1)}=\left(p_{1},\dots,p_{d_{1}}\right)^{\text{T}},\text{ and }p^{(2)}=\left(p_{d_{1}+1},\dots,p_{n}\right)^{\text{T}}.$
The partial transpose (PT) in the first partition is a transformation on the Wigner function as follows, see e. g. [Perescrit, SimonR]:
(1.12) $W\big{(}x^{(1)},p^{(1)};x^{(2)},p^{(2)}\big{)}\rightarrow\widetilde{W}(x,p)\stackrel{{\scriptstyle\text{def}}}{{=}}W\big{(}x^{(1)},-p^{(1)};x^{(2)},p^{(2)}\big{)}.$
A similar definition for the partial transpose in the second partition is equivalent for all practical purposes, see e. g. [GUHNE] for more details. In fact, this property easily follows in the position representation. The following necessary condition of separability were independently found by Peres [Perescrit] and the Horodecki family [HORODECKI19961].
Theorem 1.19 (Peres–Horodecki criterion).
If $\widehat{\rho}$ is separable, then $\widetilde{W}(x,p)$ defines a density operator.
It appears that the another direction of the above theorem also holds if $\widehat{\rho}_{G}$ is Gaussian and one set in the partition contains a single element, that is, either $d_{1}=1$ or $d_{2}=1$. The following theorem was first proved in the case $d_{1}=d_{2}=1$ by Simon [SimonR], and the proof was generalized for the $d_{1}=1$ or $d_{2}=1$ case by Werner and Wolf [WernerWolf]. Later Lami, Serafini and Adesso [Lami] simplified the proof and summarized the current status of knowledge about this topic.
Theorem 1.20 (Werner–Wolf).
Let $\widehat{\rho}_{G}$ be Gaussian and let $d_{1}=1$ or $d_{2}=1$. Then $\widehat{\rho}_{G}$ is separable if and only if $\widetilde{W}_{G}(x,p)$ represents a Gaussian density operator.
The kernel $\rho(x,y)$ in position representation can be given according to (1.7), which is separable if and only if we can write it in the following form:
(1.13) $\rho(x,y):=\rho\big{(}x^{(1)},y^{(1)};x^{(2)},y^{(2)}\big{)}=\sum_{j=1}^{\infty}\pi_{j}\rho^{(1)}_{j}\big{(}x^{(1)},y^{(1)}\big{)}\rho^{(2)}_{j}\big{(}x^{(2)},y^{(2)}\big{)},$
where $\pi_{j}\geq 0$ and $\sum_{j=1}^{\infty}\pi_{j}=1$, and $\rho^{(i)}_{j}$ is the position representation of the density operator $\widehat{\rho}^{\,(i)}_{j}$ for all $i\in\{1,2\}$ and $j\geq 1$, and we used the notation
$x$ $=(x_{1},\dots,x_{n})^{\text{T}},\leavevmode\nobreak\ x^{(1)}=\left(x_{1},\dots,x_{d_{1}}\right)^{\text{T}},\text{ and }x^{(2)}=\left(x_{d_{1}+1},\dots,x_{n}\right)^{\text{T}};$
$y$ $=(y_{1},\dots,y_{n})^{\text{T}},\leavevmode\nobreak\ y^{(1)}=\left(y_{1},\dots,y_{d_{1}}\right)^{\text{T}},\text{ and }y^{(2)}=\left(y_{d_{1}+1},\dots,y_{n}\right)^{\text{T}},$
see e. g. [Murphy, Section 6.3]. The condition $\operatorname{Tr}(\widehat{\rho})=1$ in position representation means
(1.14) $\int_{\mathbb{R}^{n}}\rho(x,x)\,\mathrm{d}x=1.$
The PT operation in position representation is given as
(1.15) $\rho(x,y)=\rho\big{(}x^{(1)},y^{(1)};x^{(2)},y^{(2)})\rightarrow\widetilde{\rho}(x,y)\stackrel{{\scriptstyle\text{def}}}{{=}}\rho\big{(}y^{(1)},x^{(1)};x^{(2)},y^{(2)}\big{)}.$
Note that if ${\widetilde{\rho}}_{2}$ denotes the PT operation in the second coordinates then
$\widetilde{\rho}(x,y)=\rho\big{(}y^{(1)},x^{(1)};x^{(2)},y^{(2)}\big{)}=\rho^{*}\big{(}x^{(1)},y^{(1)};y^{(2)},x^{(2)}\big{)}=({\widetilde{\rho}}_{2}(x,y))^{*},$
and conjugating the kernel does not change the spectrum of a self-adjoint operator, so the two possible forms of the PT transform are equivalent for our investigation. For the sake of completeness, here we enclose Theorems 1.19 and 1.20 in position representation, too.
Definition 1.21.
A density operator $\widehat{\rho}$ is called PPT (positive partial transpose), if its partial transpose $\widetilde{\rho}(x,y)$ is a kernel of a density operator. A density operator is called NPT (non-positive partial transpose) if it is not PPT.
Theorem 1.22 (Peres–Horodecki criterion).
If $\widehat{\rho}$ is a separable, then it is PPT.
Theorem 1.23 (Werner–Wolf).
Let $\widehat{\rho}_{G}$ be Gaussian and let $d_{1}=1$ or $d_{2}=1$. Then $\widehat{\rho}_{G}$ is separable if and only if it is PPT.
Definition 1.24.
Let $P$ be a self-adjoint polynomial and $\widehat{\rho}_{G}$ be a Gaussian density operator and assume that $P\rho_{G}$ is a kernel of a positive operator. Then we can define the polynomial Gaussian density operator $\widehat{\rho}_{\operatorname{PG}}$ such that its kernel $\rho_{\operatorname{PG}}$ is a constant multiple of $P\rho_{G}$, see (1.14) for the normalizing factor.
Recall the preorder $\preceq$ from Definition 1.7. The next corollary easily follows from Theorem 1.9. It means that if we have an NPT polynomial Gaussian operator, then we can find a lot of others by changing only its Gaussian factor.
Corollary 1.25.
Let $\widehat{\rho}_{G_{0}}$, $\widehat{\rho}_{G_{1}}$ be Gaussian density operators with $\widetilde{\rho}_{G_{0}}\preceq\widetilde{\rho}_{G_{1}}$. Let $P$ be a self-adjoint polynomial and assume that $\widehat{\rho}_{\operatorname{PG}_{0}}$ and $\widehat{\rho}_{\operatorname{PG}_{1}}$ exist. Then
$\widehat{\rho}_{\operatorname{PG}_{1}}\text{ is NPT}\quad\Longrightarrow\quad\widehat{\rho}_{\operatorname{PG}_{0}}\text{ is NPT}.$
From now on let $P$ be a self-adjoint polynomial and let $\widehat{\rho}_{G}$ be a Gaussian density operator such that the corresponding polynomial Gaussian density operator $\widehat{\rho}_{\operatorname{PG}}$ exists. Theorem 1.2 yields the next corollary.
Corollary 1.26.
If $\widehat{\rho}_{G}$ is NPT, then $\widehat{\rho}_{\operatorname{PG}}$ is NPT as well.
Theorems 1.2 and 1.23 imply the following corollary.
Corollary 1.27.
If $d_{1}=1$ or $d_{2}=1$ and $\widehat{\rho}_{G}$ is entangled, then $\widehat{\rho}_{\operatorname{PG}}$ is entangled.
The following problem asks whether Theorem 1.23 can be generalized to polynomial Gaussian operators, that is, in the $d_{1}=1$ or $d_{2}=1$ case the entanglement reduces to a positivity check.
Problem 1.28.
Let $d_{1}=1$ or $d_{2}=1$. If $\widehat{\rho}_{\operatorname{PG}}$ is entangled, is it necessarily NPT? In other words, is it true that $\widehat{\rho}_{\operatorname{PG}}$ is separable if and only if it is PPT?
1.4. Structure of the paper
We prove Theorem 1.2 in Section 2. First we prove it in one dimension in Subsection 2.1, then we summarize the knowledge on the symplectic decomposition of Gaussian operators in Subsection 2.2, and finally we fully prove Theorem 1.2 in Subsection 2.3 by tracing back the general case to the one-dimensional result. Section 3 is dedicated to the proof of Theorem 1.3. The results concerning our preorder will be proved in Section 4, namely Theorems 1.9 and 1.15, Claim 1.11, and Corollary 1.13.
2. When the Gaussian is not positive
The main goal of this section is to prove Theorem 1.2. First we will settle the one-dimensional case in Subsection 2.1. Then in Subsection 2.2 we recall the symplectic decomposition for Gaussian kernels, in particular we state Claim 2.4 which will be useful in Section 4, too. Finally, in Subsection 2.3 we trace back the general case to the one-dimensional result using a symplectic decomposition based on Williamson’s theorem [Williamson_origin].
2.1. The one-dimensional case
In this subsection we prove the following.
Theorem 2.1.
Assume that $P\colon\mathbb{R}^{2}\to\mathbb{C}$ is a non-zero self-adjoint polynomial and let $\kappa_{G}\colon\mathbb{R}^{2}\to\mathbb{R}$ be a Gaussian kernel of the form
$\kappa_{G}(x,y)=\exp\left(-A(x-y)^{2}-C(x+y)^{2}\right)$
such that $\widehat{\kappa}_{G}$ is not positive semidefinite, that is, $C>A>0$. Then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite.
Proof.
Assume to the contrary that $\widehat{\kappa}_{\operatorname{PG}}$ is positive semidefinite. First we show that $R(x)=P(x,x)$ is a non-zero polynomial. Indeed, assume to the contrary that $R(x)=P(x,x)=0$ for all $x\in\mathbb{R}$, then by (1.4) the sum of the eigenvalues equals to
(2.1) $\operatorname{Tr}\left(\widehat{\kappa}_{\operatorname{PG}}\right)=\int_{-\infty}^{\infty}\kappa_{\operatorname{PG}}(x,x)\,\mathrm{d}x=0.$
As $\widehat{\kappa}_{\operatorname{PG}}$ is not the zero operator, it has non-zero eigenvalues. Hence $\widehat{\kappa}_{\operatorname{PG}}$ has a negative eigenvalue by (2.1), which contradicts that it is positive definite.
Now define the polynomial $Q\colon\mathbb{R}^{2}\to\mathbb{C}$ as
$Q(x,y)=P(x,-x+y)+P(-x+y,x).$
As $Q(x,2x)=2P(x,x)$, it follows that $Q$ is not the zero polynomial. Therefore, we can fix $\varepsilon\in\mathbb{R}$ such that $Q(x,\varepsilon)$ is a non-zero polynomial of $x$. Applying Theorem 1.1 for $\kappa_{\operatorname{PG}}$ with $n=1$, $k=2$, $c_{1}=1$, and $c_{2}=-1$ implies that for all $x_{1},x_{2}\in\mathbb{R}$ we have
(2.2) $\kappa_{\operatorname{PG}}(x_{1},x_{1})+\kappa_{\operatorname{PG}}(x_{2},x_{2})\geq\kappa_{\operatorname{PG}}(x_{1},x_{2})+\kappa_{\operatorname{PG}}(x_{2},x_{1}).$
Substituting $x_{1}=x$ and $x_{2}=\varepsilon-x$ into (2.2) and using the notation
(2.3) $E(x)=\exp\left(4(A-C)x^{2}-4A\varepsilon x+(A+C)\varepsilon^{2}\right)$
we obtain that for all $x\in\mathbb{R}$ we have
(2.4) $E(x)\left[R(x)+R(\varepsilon-x)\exp(8C\varepsilon x-4C\varepsilon^{2})\right]\geq Q(x,\varepsilon).$
The term $4(A-C)x^{2}$ in (2.3) and $A-C<0$ imply that the left hand side of (2.4) tends to 00 as $x\to\infty$. As $Q(x,\varepsilon)$ is a non-zero polynomial of $x$, the limit ${\lim_{x\to\infty}Q(x,\varepsilon)\neq 0}$ exists, because $Q$ is either constant or the limit is +$\infty$+ ∞ or $\infty$- ∞. These and (2.4) imply that
(2.5) $\lim_{x\to\infty}Q(x,\varepsilon)<0.$
Now consider the kernel
$\tau(x,y)=xy\kappa_{\operatorname{PG}}(x,y).$
As $\kappa_{\operatorname{PG}}$ satisfies the Mercer-type inequalities (1.5), so does $\tau$. Therefore, similarly to (2.2) for all $x_{1},x_{2}\in\mathbb{R}$ we obtain that
(2.6) $\tau(x_{1},x_{1})+\tau(x_{2},x_{2})\geq\tau(x_{1},x_{2})+\tau(x_{2},x_{1}).$
Substituting $x_{1}=x$ and $x_{2}=\varepsilon-x$ into (2.6) yields that for all $x\in\mathbb{R}$ we have
(2.7) $E(x)\left[x^{2}R(x)+(\varepsilon-x)^{2}R(\varepsilon-x)\exp(8C\varepsilon x-4C\varepsilon^{2})\right]\geq x(\varepsilon-x)Q(x,\varepsilon).$
The term $4(A-C)x^{2}$ in (2.3) and $A-C<0$ imply that the left hand side of (2.7) tends to 00 as $x\to\infty$. Since ${\lim_{x\to\infty}x(\varepsilon-x)=-\infty}$, inequality (2.5) yields ${\lim_{x\to\infty}x(\varepsilon-x)Q(x,\varepsilon)=+\infty}$, which clearly contradicts (2.7). ∎
2.2. Symplectic decomposition and Williamson’s theorem
The results of this subsection are known, but we will need them in Subsection 2.3 and also in Section 4. Recall that our Gaussian kernel $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ is of the form
$\kappa_{G}(x,y)=\exp\left\{-(x-y)^{\text{T}}A(x-y)-i(x-y)^{\text{T}}B(x+y)-(x+y)^{\text{T}}C(x+y)\right\},$
where $A,B,C\in\mathbb{R}^{n\times n}$ such that $A$ and $C$ are positive definite and $B$ is arbitrary. It is well known that in dimension $1$ the operator $\widehat{\kappa}_{G}$ is positive semidefinite if and only if $A\geq C>0$, see e. g. [Newton] where even the eigenfunctions are calculated.
Calculating the Wigner–Weyl transform of $\kappa_{G}$ (see [SimonWW, (2.18) and (2.19)] with a slightly different terminology) yields the following formula in the phase space:
(2.8) $W_{G}(x,p)=c_{G}\exp\left\{-v^{\text{T}}Gv\right\},$
where $c_{G}=2^{n}\pi^{-\frac{3}{2}n}(\det C)^{\frac{3}{2}}(\det A)^{-\frac{1}{2}}$, and $v=(x,p)^{\text{T}}$, and
(2.9) $G=\begin{pmatrix}4C+B^{\text{T}}A^{-1}B&&\frac{1}{2}B^{\text{T}}A^{-1}\\ \frac{1}{2}A^{-1}B&&\frac{1}{4}A^{-1}\end{pmatrix}$
is a symmetric, positive definite [^5] For the positivity of $G$ one can simply check that $v^{\text{T}}Gv\geq 0$ for all $v\in\mathbb{R}^{2n}$; it also follows from the fact that the Wigner–Weyl transform maps $L^{2}(\mathbb{R}^{2n})$ to $L^{2}(\mathbb{R}^{2n})$, so $W$ is square integrable. real $2n\times 2n$ matrix [^6] In the literature $G^{-1}$ is called the Gaussian covariance matrix.. Define the $2n\times 2n$ matrix
$\Omega=\begin{pmatrix}0&I_{n}\\ -I_{n}&0\end{pmatrix},$
where $I_{n}$ is the $n\times n$ identity matrix and 00 is the $n\times n$ zero matrix. A $2n\times 2n$ real matrix $S\in\operatorname{Sp}(2n,\mathbb{R})$ is called symplectic if it satisfies $S^{\text{T}}\Omega S=\Omega$.
For the following important theorem see [Gosson_Harmonic, Theorem 215], which states that applying a symplectic transform in the phase space does not change the spectrum of the operator.
Theorem 2.2 (Weyl correspondence).
Assume that $\kappa\in L^{2}(\mathbb{R}^{2n})$ and $W$ is its Wigner–Weyl transform, and let $S\in\operatorname{Sp}(2n,\mathbb{R})$ be a symplectic matrix. Then the operator corresponding to $W\circ S$ equals $\widehat{S}^{-1}\widehat{\kappa}\widehat{S}$ with some operator [^7] Note that the operator $\widehat{S}$ is explicitly constructed in [deGosson_book, Serafini, _pereira]. $\widehat{S}$; note that $\widehat{\kappa}$ and $\widehat{S}^{-1}\widehat{\kappa}\widehat{S}$ are in the same conjugacy class.
The following theorem is due to Williamson [Williamson_origin], see also [Serafini, Subsection 3.2.3], [Gosson_Harmonic, Theorem 93], or [deGosson_book, Theorem 8.11].
Theorem 2.3 (Williamson’s theorem).
Let $G\in\mathbb{R}^{2n\times 2n}$ be a symmetric, positive definite matrix. There is a $2n\times 2n$ symplectic matrix [^8] Note that finding such a matrix $S$ is not easy in practice, see [_pereira] for tackling this problem. $S\in\operatorname{Sp}(2n,\mathbb{R})$ such that
(2.10) $S^{\text{T}}GS=\begin{pmatrix}\Lambda&0\\ 0&\Lambda\end{pmatrix},$
where $\Lambda$ is the $n\times n$ diagonal matrix
$\Lambda=\begin{bmatrix}\mu_{1}&&\\ &\ddots&\\ &&\mu_{n}\end{bmatrix}.$
Furthermore, $\mu_{k}$ are positive and ±$\pm\mu_{k}$ are the eigenvalues [^9] The reciprocals $1/\mu_{i}$ are called the symplectic eigenvalues of the Gaussian covariance matrix. of the matrix $M=iG\Omega$.
By Theorem 2.3 we can diagonalize $G$ from (2.9) with a symplectic matrix, that is, there is a symplectic matrix $S\in\operatorname{Sp}(2n,\mathbb{R})$ such that (2.10) holds. We apply the linear transformation $(x^{\prime},p^{\prime})^{\text{T}}=v^{\prime}=Sv$ in the phase space; by Theorem 2.2 this does not change the spectrum of $\widehat{\kappa}_{G}$. Now (2.8) and (2.10) imply that the linear transformation $v\mapsto Sv$ yields the product formula
$W_{G}(x^{\prime},p^{\prime})=c_{G}\prod^{n}_{k=1}W_{G_{k}}(x^{\prime}_{k},p^{\prime}_{k}),$
where
(2.11) $W_{G_{k}}(x,p)=\exp\left\{-\mu_{k}(x^{2}+p^{2})\right\}.$
The inverse Wigner transformation (1.7) allows us to factorize our kernel in position representation as
(2.12) $\kappa^{\prime}_{G}(x^{\prime},y^{\prime})=c_{G}\prod^{n}_{k=1}\kappa_{G_{k}}(x^{\prime}_{k},y^{\prime}_{k}),$
where (2.11) yields
(2.13) $\kappa_{G_{k}}(x,y)=\left(\frac{\pi}{\mu_{k}}\right)^{\frac{1}{2}}\exp\left\{-\frac{1}{4\mu_{k}}(x-y)^{2}-\frac{\mu_{k}}{4}(x+y)^{2}\right\}.$
Clearly, $\widehat{\kappa}^{\prime}_{G}\geq 0$ if and only if $\widehat{\kappa}_{G_{k}}\geq 0$ for all $k$, and by (2.13) this is equivalent to $\mu_{k}\leq 1$ for all $1\leq k\leq n$. We emphasize this result in the following claim.
Claim 2.4 (Positivity of Gaussian operators).
The Gaussian operator $\widehat{\kappa}_{G}$ is positive semidefinite if and only if all the eigenvalues $\mu_{k}$ from Theorem 2.3 satisfy $\mu_{k}\leq 1$.
2.3. The general case
Before proving Theorem 1.2 we recall the following notion tailor-made for our need. Define the partial trace [^10] In physics the partial trace operation is applied for kernels of density operators. (see [Partial_trace]) of a kernel $\kappa\in\mathcal{S}(\mathbb{R}^{2n})$ in the coordinates $x_{2},\dots,x_{n}$ as the kernel $\eta\in\mathcal{S}(\mathbb{R}^{2})$ satisfying
$\eta(x_{1},y_{1})=\int_{\mathbb{R}^{n-1}}\kappa(x_{1},x_{2},\dots,x_{n},y_{1},x_{2},\dots,x_{n})\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}.$
Fact 2.5.
$\widehat{\kappa}$ and $\widehat{\eta}$ are trace class with $\operatorname{Tr}(\widehat{\eta})=\operatorname{Tr}(\widehat{\kappa})$, and $\widehat{\kappa}\geq 0$ implies $\widehat{\eta}\geq 0$.
Proof.
As $\kappa\in\mathcal{S}(\mathbb{R}^{2n})$ implies $\eta\in\mathcal{S}(\mathbb{R}^{2})$, we obtain that $\widehat{\kappa}$ and $\widehat{\eta}$ are trace class. Applying (1.4) for both $\widehat{\eta}$ and $\widehat{\kappa}$, and using the definition of $\eta$ imply
$\operatorname{Tr}(\widehat{\eta})=\int_{\mathbb{R}}\eta(x_{1},x_{1})\,\mathrm{d}x_{1}=\int_{\mathbb{R}^{n}}\kappa(x,x)\,\mathrm{d}x=\operatorname{Tr}(\widehat{\kappa}).$
Now assume that $\widehat{\kappa}\geq 0$. Let $x_{1},\dots,x_{k}\in\mathbb{R}$ and $c_{1},\dots,c_{k}\in\mathbb{C}$ be arbitrary. Theorem 1.1 implies that for every $u\in\mathbb{R}^{n-1}$ we have
(2.14) $\sum_{i,j=1}^{k}c_{i}c^{*}_{j}\kappa(x_{i},u;x_{j},u)\geq 0.$
The definition of $\eta$ and integrating both sides of (2.14) with respect to $u$ imply
$\sum_{i,j=1}^{k}c_{i}c^{*}_{j}\eta(x_{i},x_{j})=\sum_{i,j=1}^{k}c_{i}c^{*}_{j}\int_{\mathbb{R}^{n-1}}\kappa(x_{i},u;x_{j},u)\,\mathrm{d}u\geq 0.$
As $x_{1},\dots,x_{k}\in\mathbb{R}$ and $c_{1},\dots,c_{k}\in\mathbb{C}$ were arbitrary, Theorem 1.1 yields $\widehat{\eta}\geq 0$. ∎
Now we are ready to prove Theorem 1.2.
Theorem 1.2.
Let $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a non-zero polynomial and let $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a Gaussian kernel. If $\widehat{\kappa}_{G}$ is not positive semidefinite, then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite, either.
Proof.
Assume to the contrary that $\widehat{\kappa}_{G}$ is not positive semidefinite but $\widehat{\kappa}_{\operatorname{PG}}$ is positive semidefinite. Recall that $\kappa_{G}$ in coordinate representation is
$\kappa_{G}(x,y)=\exp\left\{-(x-y)^{\text{T}}A(x-y)-i(x-y)^{\text{T}}B(x+y)-(x+y)^{\text{T}}C(x+y)\right\},$
where $A,B,C\in\mathbb{R}^{n\times n}$, and $A,C$ are symmetric and positive definite. We will use the notation and results of Subsection 2.2. We have seen in (2.8) that the Wigner–Weyl transform of $\kappa_{G}$ is
$W_{G}(x,p)=c_{G}\exp\left\{-v^{\text{T}}Gv\right\},$
where $v=(x,p)^{\text{T}}$ and the constant $c_{G}>0$ and the symmetric, positive definite matrix $G\in\mathbb{R}^{2n\times 2n}$ are given there. By differentiation under the integral sign we obtain that the Wigner–Weyl transform of $\kappa_{\operatorname{PG}}$ satisfies
(2.15) $W_{\operatorname{PG}}(x,p)=Q(x,p)\exp\left\{-v^{\text{T}}Gv\right\},$
where $Q$ is a polynomial in $2n$ variables. By Theorem 2.3 we can diagonalize $G$ by a symplectic matrix $S\in\operatorname{Sp}(2n,\mathbb{R})$, that is,
(2.16) $S^{\text{T}}GS=\begin{pmatrix}\Lambda&0\\ 0&\Lambda\end{pmatrix},$
where $\Lambda\in\mathbb{R}^{n\times n}$ is a diagonal matrix with positive diagonal entries $\mu_{k}$.
We will apply the linear transformation $(x^{\prime},p^{\prime})^{\text{T}}=v^{\prime}=Sv$ in the phase space. By Theorem 2.2 the operator corresponding to $W_{\operatorname{PG}}\circ S$ has the same spectrum as $\widehat{\kappa}_{\operatorname{PG}}$, so it remains positive semidefinite. Now (2.15) and (2.16) imply that the linear transformation $v\mapsto Sv$ yields the product formula
$W_{\operatorname{PG}}(x^{\prime},p^{\prime})=Q(x,p)\prod^{n}_{k=1}W_{G_{k}}(x^{\prime}_{k},p^{\prime}_{k}),$
where
$W_{G_{k}}(x,p)=\exp\left\{-\mu_{k}(x^{2}+p^{2})\right\}.$
We have already calculated the inverse Wigner–Weyl transform (1.7) of the Gaussian part $\prod_{k=1}^{n}W_{G_{k}}$, see (2.8). By differentiating under the integral sign we easily obtain that our inverse Wigner–Weyl transform only changes by a polynomial factor, that is, there is a polynomial $R$ in $2n$ variables such that our kernel in position representation is
$\kappa^{\prime}_{\operatorname{PG}}(x^{\prime},y^{\prime})=R(x^{\prime},y^{\prime})\prod^{n}_{k=1}\kappa_{G_{k}}(x^{\prime}_{k},y^{\prime}_{k}),$
where $\kappa_{G_{k}}$ are the one-dimensional Gaussian kernels calculated in (2.13). As $\widehat{\kappa}_{G}$ is not positive semidefinite, and in $\kappa^{\prime}_{G}$ the same factors $\kappa_{G_{k}}$ appear, we obtain that there exists $k$ such that $\widehat{\kappa}_{G_{k}}$ is not positive semidefinite. We may assume without loss of generality that this $k$ equals $1$. Now we can define the two-variable polynomial
$S(x^{\prime}_{1},y^{\prime}_{1})=\int_{\mathbb{R}^{n-1}}R(x^{\prime}_{1},\dots,x^{\prime}_{n},y^{\prime}_{1},x^{\prime}_{2},\dots,x_{n}^{\prime})\prod^{n}_{k=2}\kappa_{G_{k}}(x^{\prime}_{k},x^{\prime}_{k})\,\mathrm{d}x^{\prime}_{2}\cdots\,\mathrm{d}x^{\prime}_{n}.$
Consider the partial trace of $\kappa^{\prime}_{\operatorname{PG}}$ in the coordinates $x^{\prime}_{2},\dots,x^{\prime}_{n}$, which yields the one-dimensional kernel $\eta$ given by
$\eta(x_{1}^{\prime},y_{1}^{\prime})=S(x_{1}^{\prime},y_{1}^{\prime})\kappa_{G_{1}}(x_{1}^{\prime},y_{1}^{\prime}).$
By Fact 2.5 the partial trace operator preserves positivity and the trace as well, so $\widehat{\eta}\geq 0$ and $S$ is not the zero polynomial. However, $\eta$ is the product of $S$ and a one-dimensional Gaussian kernel $\kappa_{G_{1}}$ such that $\widehat{\kappa}_{G_{1}}$ is not positive semidefinite, so the positivity of $\widehat{\eta}$ contradicts the one-dimensional result Theorem 2.1. ∎
3. When the polynomial is of odd degree
The main goal of this section is to prove Theorem 1.3, but first we need some preparation. We will need the Fourier transformation $\mathcal{F}\colon L^{2}(\mathbb{R}^{2n})\to L^{2}(\mathbb{R}^{2n})$, here we use the convention that for all $f\in L^{2}(\mathbb{R}^{2n})$ and $u,v\in\mathbb{R}^{n}$ we have
(3.1) $\mathcal{F}(f)(u,v)=\iint_{\mathbb{R}^{2n}}f(x,y)e^{-i\left(u^{\text{T}}x+v^{\text{T}}y\right)}\,\mathrm{d}x\,\mathrm{d}y.$
We define the integral transformation $\mathcal{L}$ for polynomial Gaussian kernels $\kappa_{\operatorname{PG}}$ such that for all $u,v\in\mathbb{C}^{n}$ we have
(3.2) $\mathcal{L}(\kappa_{\operatorname{PG}})(u,v)=\iint_{\mathbb{R}^{2n}}\kappa_{\operatorname{PG}}(x,y)\exp\left(-iu^{\text{T}}(x-y)-v^{\text{T}}(x+y)\right)\,\mathrm{d}x\,\mathrm{d}y.$
In the following lemma we calculate the $\mathcal{L}$-transform of a Gaussian kernel.
Lemma 3.1.
There is a positive constant $c=c(A,B,C)$ such that for all $u,v\in\mathbb{C}^{n}$ our integral $\mathcal{L}(\kappa_{G})(u,v)$ equals to
$c\cdot\exp\left[-\left(\frac{1}{2}BC^{-1}v-u\right)^{\text{T}}\left(4A+BC^{-1}B^{\text{T}}\right)^{-1}\left(\frac{1}{2}BC^{-1}v-u\right)+\frac{1}{4}v^{\text{T}}C^{-1}v\right].$
Proof.
For a positive definite real matrix $M$ and $b\in\mathbb{C}^{n}$ the following formula is well known, see e. g. [Zee2010, Section I.2. Appendix 2] for the proof:
(3.3) $\int_{\mathbb{R}^{n}}\exp\left[-x^{\text{T}}Mx+b^{\text{T}}x\right]\,\mathrm{d}x=\frac{\pi^{\frac{n}{2}}}{\sqrt{\det M}}\exp{\left(\frac{1}{4}b^{\text{T}}M^{-1}b\right)}.$
First we make a change of variables from $x,y$ to $r=x-y$ and $R=\frac{x+y}{2}$, note that the determinant of its Jacobian matrix has absolute value $1$. Then using (3.3) with $M_{1}=4C$ and $b_{1}=-2(v+iB^{\text{T}}r)$, and again with $M_{2}=A+\frac{1}{4}BC^{-1}B^{\text{T}}$ and $b_{2}=i\left(\frac{1}{2}BC^{-1}v-u\right)$ we obtain that
$\mathcal{L}(\kappa_{G})(u,v)=\iint_{\mathbb{R}^{2n}}\exp\left[-r^{\text{T}}Ar-2ir^{\text{T}}BR-4R^{\text{T}}CR-iu^{\text{T}}r-2v^{\text{T}}R\right]\,\mathrm{d}R\,\mathrm{d}r$
$=\frac{\pi^{\frac{n}{2}}\exp\left(\frac{1}{4}v^{\text{T}}C^{-1}v\right)}{\sqrt{\det(4C)}}\int_{\mathbb{R}^{n}}\exp\left[-r^{\text{T}}\left(A+\frac{1}{4}BC^{-1}B^{\text{T}}\right)r+i\left(\frac{1}{2}v^{\text{T}}C^{-1}v\right)r\right]\,\mathrm{d}r$
$=c\exp\left[-\left(\frac{1}{2}BC^{-1}v-u\right)^{\text{T}}\left(4A+BC^{-1}B^{\text{T}}\right)^{-1}\left(\frac{1}{2}BC^{-1}v-u\right)+\frac{1}{4}v^{\text{T}}C^{-1}v\right],$
where $c=\pi^{n}\left(\det\left(4AC+BC^{-1}B^{\text{T}}C\right)\right)^{-\frac{1}{2}}$. This finishes the proof. ∎
Before proving Theorem 1.3 we also need a couple of facts.
Fact 3.2.
Let $Q\colon\mathbb{R}^{2n}\to\mathbb{C}$ be a polynomial with $2n$ real variables.
  1. (i)
    If $Q$ takes only real values, then all of its coefficients are real;
  2. (ii)
    If $Q$ has real coefficients and $\deg Q$ is odd, then $Q$ takes negative values.
Proof.
First we prove (i). We can write $Q=Q_{1}+iQ_{2}$, where the polynomials $Q_{1},Q_{2}\colon\mathbb{R}^{2n}\to\mathbb{R}$ have real coefficients. As $Q$ takes only real values, we obtain that $Q_{2}(x,y)=0$ for all $x,y\in\mathbb{R}^{n}$, which implies that $Q_{2}\equiv 0$. Then $Q=Q_{1}$, and the claim follows.
Now we prove (ii). Let $d=\deg Q$, which is odd by our assumption. Let us decompose $Q$ as $Q=Q_{1}+Q_{2}$ such that $Q_{1},Q_{2}\colon\mathbb{R}^{2n}\to\mathbb{R}$ are polynomials with real coefficients, and $Q_{1}$ contains all the monomials of $Q$ with degree $d$. Then clearly $\deg Q_{2}<d$ and $Q_{1}$ is homogeneous of degree $d$, that is, $Q_{1}(cx,cy)=c^{d}Q_{1}(x,y)$ for all $c\in\mathbb{R}$ and $x,y\in\mathbb{R}^{n}$. Fix $x_{0},y_{0}\in\mathbb{R}^{n}$ such that $Q_{1}(x_{0},y_{0})\neq 0$, and let $r_{0}=Q_{1}(x_{0},y_{0})$. Let us define the one-variable polynomial $R\colon\mathbb{R}\to\mathbb{R}$ as
$R(c)=Q(cx_{0},cy_{0})=r_{0}c^{d}+Q_{2}(cx_{0},cy_{0}).$
As $\deg Q_{2}<d$, we have $\deg R=d$ and its leading coefficient is $r_{0}\neq 0$. As $d$ is odd, $R$ takes negative values, which implies that $Q$ takes negative values as well. ∎
Fact 3.3.
Assume that $\kappa\in L^{2}(\mathbb{R}^{2n})$ is a self-adjoint kernel such that $\widehat{\kappa}$ is positive semidefinite. Let $g\colon\mathbb{R}^{n}\to\mathbb{C}$ be a continuous function and suppose that the function $\tau(x,y)=g(x)g^{*}(y)\kappa(x,y)$ satisfies $\tau\in L^{1}(\mathbb{R}^{2n})$. Then $\iint_{\mathbb{R}^{2n}}\tau(x,y)\,\mathrm{d}x\,\mathrm{d}y\geq 0$.
Proof.
As $\tau\in L^{1}(\mathbb{R}^{2n})$, the integral $\iint_{\mathbb{R}^{2n}}\tau(x,y)\,\mathrm{d}x\,\mathrm{d}y$ exists. For positive integers $n$ define $g_{n}\colon\mathbb{R}^{n}\to\mathbb{C}$ such that $g_{n}(x)=g(x)$ if $|x|\leq n$ and $g_{n}(x)=0$ otherwise. As $g$ is continuous, we have $g_{n}\in L^{2}(\mathbb{R}^{n})$. Therefore, the positivity of $\widehat{\kappa}$ implies that for all $n$ we have
$0\leq\iint_{\mathbb{R}^{2n}}g_{n}(x)g_{n}^{*}(y)\kappa(x,y)\,\mathrm{d}x\,\mathrm{d}y\to\iint_{\mathbb{R}^{2n}}\tau(x,y)\,\mathrm{d}x\,\mathrm{d}y$
as $n\to\infty$, hence $\iint_{\mathbb{R}^{2n}}\tau(x,y)\,\mathrm{d}x\,\mathrm{d}y\geq 0$ holds. ∎
Now we are able to prove Theorem 1.3.
Theorem 1.3.
Assume that $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ is a polynomial of odd degree, and $\kappa_{G}\colon\mathbb{R}^{2n}\to\mathbb{C}$ is any Gaussian kernel. Then $\widehat{\kappa}_{\operatorname{PG}}$ is not positive semidefinite.
Proof.
Assume to the contrary that $\widehat{\kappa}_{\operatorname{PG}}$ is positive semidefinite. Recall the integral transformation $\mathcal{L}$ from (3.2), and note that
$\exp\left(-iu^{\text{T}}(x-y)-v^{\text{T}}(x+y)\right)=g_{u,v}(x)g_{u,v}^{*}(y),$
where $g_{u,v}(x)=\exp(-iu^{\text{T}}x-v^{\text{T}}x)$. Thus Fact 3.3 implies that
(3.4) $\mathcal{L}(\kappa_{\operatorname{PG}})(u,v)\geq 0\quad\text{for all}\quad u,v\in\mathbb{R}^{n}.$
We want to calculate $\mathcal{L}(\kappa_{\operatorname{PG}})$ by differentiating under the integral sign of $\mathcal{L}(\kappa_{G})$. We define the differential operators $D_{i}$ for all $1\leq i\leq 2n$ such that.
$D_{i}=\begin{cases}\frac{1}{2}\left(i\frac{\partial}{\partial u_{i}}-\frac{\partial}{\partial v_{i}}\right)&\textrm{if }1\leq i\leq n\\ -\frac{1}{2}\left(i\frac{\partial}{\partial u_{i}}+\frac{\partial}{\partial v_{i}}\right)&\textrm{if }n+1\leq i\leq 2n.\end{cases}$
For any polynomial $R=R(x,y)$ and $1\leq i\leq n$ by Lemma 3.1 we obtain that
(3.5) $D_{i}\mathcal{L}(R\kappa_{G})=\mathcal{L}(x_{i}R\kappa_{G})\quad\text{and}\quad D_{n+i}\mathcal{L}(R\kappa_{G})=\mathcal{L}(y_{i}R\kappa_{G}).$
For the polynomial $P$ choose the set $\mathcal{I}\subset\mathbb{N}^{2n}$ such that
$P(x,y)=\sum_{(k_{1},\dots,k_{2n})\in\mathcal{I}}a(k_{1},\dots,k_{2n})x_{1}^{k_{1}}\cdots x_{n}^{k_{n}}y_{1}^{k_{n+1}}\cdots y_{n}^{k_{2n}}.$
Let us define the differential operator $P(D)$ as
$P(D)=\sum_{(k_{1},\dots,k_{2n})\in\mathcal{I}}a(k_{1},\dots,k_{2n})D_{1}^{k_{1}}\cdots D_{2n}^{k_{2n}}.$
Then (3.5) implies that
(3.6) $\mathcal{L}(\kappa_{\operatorname{PG}})=P(D)\mathcal{L}(\kappa_{G}).$
Clearly $(D_{1}^{k_{1}}\cdots D_{2n}^{k_{2n}})\mathcal{L}(\kappa_{G})=Q_{k_{1}\dots k_{2n}}\mathcal{L}(\kappa_{G})$, where $Q_{k_{1}\dots k_{2n}}$ is a polynomial of degree exactly $k_{1}+\dots+k_{2n}$. Therefore, by (3.6) we obtain that there is a polynomial $Q$ such that $\deg Q\leq\deg P$ and
(3.7) $\mathcal{L}(\kappa_{\operatorname{PG}})=Q\mathcal{L}(\kappa_{G}).$
As $\mathcal{L}(\kappa_{\operatorname{PG}})$ takes only real values, so does $Q$, hence the coefficients of $Q$ are reals by Fact 3.2 (i). Now it is enough to show that $\deg Q=\deg P$. Indeed, assume that $\deg Q=\deg P$. Then by Fact 3.2 (ii) the odd degree polynomial $Q$ takes negative values, so $\mathcal{L}(\kappa_{\operatorname{PG}})$ takes negative values by (3.7) as well, but this contradicts (3.4). Note that it is unclear yet why $\deg Q<\deg P$ cannot happen due to cancellations in the linear combination of the polynomials $Q_{k_{1}\dots k_{2n}}$, $(k_{1},\dots,k_{2n})\in\mathcal{I}$.
Finally, we prove $\deg Q=\deg P$. Assume to the contrary that $\deg Q<\deg P$. Observe that (3.2) provides an easy connection between the transforms $\mathcal{F}$ and $\mathcal{L}$. Applying this twice with (3.7) implies that
(3.8) $\begin{split}\mathcal{F}(\kappa_{\operatorname{PG}})(u,v)&=\mathcal{L}(\kappa_{\operatorname{PG}})\left(\frac{1}{2}(u+v),\frac{i}{2}(v-u)\right)\\ &=Q\left(\frac{1}{2}(u+v),\frac{i}{2}(v-u)\right)\mathcal{L}(\kappa_{G})\left(\frac{1}{2}(u+v),\frac{i}{2}(v-u)\right)\\ &=Q^{\prime}(u,v)\mathcal{F}(\kappa_{G})(u,v),\end{split}$
where $Q^{\prime}(u,v)=Q((u+v)/2,i(v-u)/2)$, so $Q^{\prime}$ is a polynomial with $\deg Q^{\prime}\leq\deg Q$. Now we will take the Fourier transform of both sides of the (3.8). On one hand, by the Fourier inversion formula we have an absolute constant $c_{n}>0$ such that
(3.9) $\mathcal{F}(\mathcal{F}(\kappa_{\operatorname{PG}}))(x,y)=c_{n}\kappa_{\operatorname{PG}}(-x,-y)=c_{n}P(-x,-y)\kappa_{G}(-x,-y).$
On the other hand, differentiating under the integral sign similarly as earlier and using the above Fourier inversion formula, we obtain that there is a polynomial $Q^{\prime\prime}$ with $\deg Q^{\prime\prime}\leq\deg Q^{\prime}$ satisfying
(3.10) $\mathcal{F}(Q^{\prime}\mathcal{F}(\kappa_{G}))(x,y)=Q^{\prime\prime}(x,y)\mathcal{F}(\mathcal{F}(\kappa_{G}))(x,y)=c_{n}Q^{\prime\prime}(x,y)\kappa_{G}(-x,-y).$
By (3.8) the Fourier transforms in (3.9) and (3.10) are equal, which yields that $Q^{\prime\prime}(x,y)=P(-x,-y)$ for all $x,y\in\mathbb{R}^{n}$. However, $\deg Q^{\prime\prime}\leq\deg Q<\deg P$ by our indirect hypothesis, which is clearly a contradiction. This completes the proof. ∎
4. A useful preorder on Gaussian kernels
One of the main goals of this section is to prove Theorem 1.9, but first we need some preparation.
Fact 4.1.
Let $\mu,\nu\in L^{2}(\mathbb{R}^{2n})$ be self-adjoint kernels such that $\widehat{\mu}$ and $\widehat{\nu}$ are positive semidefinite. Then $\mu\nu\in L^{2}(\mathbb{R}^{2n})$ and $\widehat{\mu\nu}$ is positive semidefinite, too.
Proof.
Clearly $\mu\nu\in L^{2}(\mathbb{R}^{2n})$. By the spectral theorem [Stein, Theorem 6.2] the operator $\widehat{\mu}$ has nonnegative eigenvalues $\{s_{i}\}_{i\geq 0}$ with eigenfunctions $\{f_{i}\}_{i\geq 0}$, and similarly $\widehat{\nu}$ has nonnegative eigenvalues $\{t_{j}\}_{j\geq 0}$ with eigenfunctions $\{g_{j}\}_{j\geq 0}$. This easily implies that the kernels $\mu,\nu$ can be written as
$\mu(x,y)=\sum_{i\geq 0}s_{i}f_{i}(x)f_{i}^{*}(y)\quad\text{and}\quad\nu(x,y)=\sum_{j\geq 0}t_{j}g_{j}(x)g_{j}^{*}(y),$
where the sums converge in $L^{2}(\mathbb{R}^{2n})$. Then clearly
$(\mu\nu)(x,y)=\sum_{k=0}^{\infty}\sum_{i+j=k}s_{i}t_{j}(f_{i}g_{j})(x)(f_{i}g_{j})^{*}(y),$
where the sum converges in $L^{2}(\mathbb{R}^{2n})$. As the kernels $(x,y)\mapsto(f_{i}g_{j})(x)(f_{i}g_{j})^{*}(y)$ define positive semidefinite operators and $s_{i}t_{j}\geq 0$ for all $i,j\geq 0$, we obtain that $\widehat{\mu\nu}$ is positive semidefinite. ∎
Fact 4.2.
Let $\tau\in L^{2}(\mathbb{R}^{2n})$ be a kernel, and let $g\colon\mathbb{R}^{n}\to\mathbb{C}$ be a Lebesgue measurable function. Define $\kappa\colon\mathbb{R}^{2n}\to\mathbb{C}$ as
$\kappa(x,y)=\tau(x,y)g(x)g^{*}(y).$
If $\widehat{\tau}$ is positive semidefinite and $\kappa\in L^{2}(\mathbb{R}^{2n})$, then $\widehat{\kappa}$ is also positive semidefinite.
Proof.
Define $h\colon\mathbb{R}^{n}\to\mathbb{R}$ as $h(x)=\exp(-|g(x)|)$. Then $h$ is Lebesgue measurable, and $0<|h(x)|\leq 1$ for all $x\in\mathbb{R}^{n}$. Let
=$\mathcal{F}=\left\{P(x)h(x)\exp(-|x|^{2}):\text{ where $P\colon\mathbb{R}^{n}\to\mathbb{C}$ is a polynomial}\right\},$
then [Schmudgen, Corollary 14.24 or 14.7 Exercise 10] yields that $\mathcal{F}$ is dense in $L^{2}(\mathbb{R}^{n})$.
Now we show that $\langle\widehat{\kappa}f,f\rangle\geq 0$ for all $f\in\mathcal{F}$. Fix an arbitrary $f\in\mathcal{F}$. As $hg$ is bounded, clearly $fg\in L^{2}(\mathbb{R}^{n})$. Then $\widehat{\tau}\geq 0$ yields $\langle\widehat{\kappa}f,f\rangle=\langle\widehat{\tau}(fg),fg\rangle\geq 0$.
Finally, we prove that $\langle\widehat{\kappa}f,f\rangle\geq 0$ for all $f\in L^{2}(\mathbb{R}^{n})$. Since $\mathcal{F}$ is dense in $L^{2}(\mathbb{R}^{n})$, there is a sequence $f_{i}\in\mathcal{F}$ such that $f_{i}\to f$ in $L^{2}(\mathbb{R}^{n})$. Clearly $\langle\widehat{\kappa}f_{i},f_{i}\rangle\to\langle\widehat{\kappa}f,f\rangle$, and we already showed that $\langle\widehat{\kappa}f_{i},f_{i}\rangle\geq 0$, thus we obtain $\langle\widehat{\kappa}f,f\rangle\geq 0$. ∎
Now we are ready to prove Theorem 1.9.
Theorem 1.9.
For $(A_{0},B_{0},C_{0}),(A_{1},B_{1},C_{1})\in\mathcal{G}(n)$ the following are equivalent:
  1. (1)
    $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$;
  2. (2)
    for each self-adjoint $\sigma\colon\mathbb{R}^{2n}\to\mathbb{C}$ if $\sigma\theta(A_{i},B_{i},C_{i})\in L^{2}(\mathbb{R}^{2n})$ for $i\in\{0,1\}$ and $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0$, then $\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0$;
  3. (3)
    for every self-adjoint Gaussian kernel $\sigma$ if $\sigma\theta(A_{i},B_{i},C_{i})\in L^{2}(\mathbb{R}^{2n})$ for $i\in\{0,1\}$ and $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})\geq 0$, then $\widehat{\sigma\theta}(A_{1},B_{1},C_{1})\geq 0$.
Proof.
[…]
(4.1) $\sigma=\theta(-A_{0}+rI_{n},-B_{0},-C_{0}+rI_{n})\in L^{2}(\mathbb{R}^{2n}).$
Then $\sigma\theta(A_{0},B_{0},C_{0})=\theta(rI_{n},0,rI_{n})\in L^{2}(\mathbb{R}^{2n})$, so clearly $\widehat{\sigma\theta}(A_{0},B_{0},C_{0})$ is positive semidefinite. Furthermore, (4.1) implies that
$\sigma\theta(A_{1},B_{1},C_{1})=\theta(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})\in L^{2}(\mathbb{R}^{2n}).$
Therefore, (3) yields that
$\widehat{\sigma\theta}(A_{1},B_{1},C_{1})=\widehat{\theta}(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})\geq 0,$
which implies (1) by definition.
[…]
$\widehat{\theta}(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})\geq 0.$
Let us define the kernels $\mu,\nu,\kappa\in L^{2}(\mathbb{R}^{2n})$ as
$\mu$ $=\sigma\theta(A_{0},B_{0},C_{0}),$
$\nu$ $=\theta(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n}),$
$\kappa$ $=\sigma\theta(A_{1},B_{1},C_{1}).$
Then we have
(4.2) $\kappa(x,y)=(\mu\nu)(x,y)g(x)g^{*}(y),$
where
$g(x)=\exp\left(r|x|^{2}\right).$
We need to prove that $\widehat{\kappa}\geq 0$. As $\widehat{\mu}\geq 0$ and $\widehat{\nu}\geq 0$, Fact 4.1 implies that $\widehat{\mu\nu}\geq 0$. Then (4.2) and Fact 4.2 with $\tau=\mu\nu$ yield that $\widehat{\kappa}\geq 0$. This completes the proof. ∎
Now we are ready to prove Claim 1.11.
Claim 1.11.
We have $(I_{n},0,I_{n})\preceq(A,B,C)$ for all $(A,B,C)\in\mathcal{G}^{+}(n)$. Hence for a self-adjoint polynomial $P\colon\mathbb{R}^{2n}\to\mathbb{C}$ the following are equivalent:
  1. (1)
    $\widehat{P\theta}(A,B,C)\geq 0$ for all $(A,B,C)\in\mathcal{G}^{+}(n)$;
  2. (2)
    $\widehat{P\theta}(I_{n},0,I_{n})\geq 0$;
  3. (3)
    $\sum_{i,j=1}^{k}c_{i}c^{*}_{j}P(x_{i},x_{j})\geq 0$ for all $x_{1},\dots,x_{k}\in\mathbb{R}^{n}$ and $c_{1},\dots,c_{k}\in\mathbb{C}$.
Proof.
First we show that $(I_{n},0,I_{n})\preceq(A,B,C)$ for all $(A,B,C)\in\mathcal{G}^{+}(n)$. Indeed, let $(A,B,C)\in\mathcal{G}^{+}(n)$ and $r=1$. Then
$(A-I_{n}+rI_{n},B,C-I_{n}+rI_{n})=(A,B,C)\in\mathcal{G}^{+}(n),$
which implies $(I_{n},0,I_{n})\preceq(A,B,C)$ by definition.
[…]
Notation 4.3.
For $U,V\in\mathbb{R}^{n\times n}$ we write $U>0$ if $U$ is positive definite, $U\geq 0$ if $U$ is positive semidefinite, and $U\geq V$ if $U-V\geq 0$.
Claim 4.4.
Let $A,B,C\in\mathbb{R}^{n\times n}$ such that $A,C$ are positive definite.
  1. (1)
    If $(A,B,C)\in\mathcal{G}^{+}(n)$ then $A\geq C$;
  2. (2)
    if $A\geq C$ and $B$ is symmetric then $(A,B,C)\in\mathcal{G}^{+}(n)$.
Proof.
First we prove (1). Applying Theorem 1.1 for $\theta=\theta(A,B,C)$ with $k=2$ and $c_{1}=1,c_{2}=-1$ for vectors $x$ and $-x$ we obtain
$\theta(x,x)+\theta(-x,-x)\geq\theta(x,-x)+\theta(-x,x),$
which easily implies that
$2\exp\big{(}-4x^{\text{T}}Cx\big{)}\geq 2\exp\big{(}-4x^{\text{T}}Ax\big{)}.$
Thus for all $x\in\mathbb{R}^{n}$ we have $x^{\text{T}}(A-C)x\geq 0$, so $A-C\geq 0$, that is, $A\geq C$.
Now we prove (2). Note that $\theta(A,B,C)(x,y)$ only differs from $\theta(A,0,C)(x,y)$ by factors of the form $f_{i,j}(x)f_{i,j}^{*}(y)$, where
$f_{i,j}(x)=\exp(-2iB_{ij}x_{i}x_{j})\quad\text{for all }1\leq i<j\leq n.$
Since these factors do not change the spectrum, we obtain that $\theta(A,B,C)\in\mathcal{G}^{+}(n)$ if and only if $\theta(A,0,C)\in\mathcal{G}^{+}(n)$. Hence we need to prove that $\theta(A,0,C)\in\mathcal{G}^{+}(n)$. By Claim 2.4 this is equivalent to the fact that all positive eigenvalues of $M=iG\Omega$ are at most $1$. We can calculate the characteristic polynomial as
$P(\lambda)=\det\left(M-\lambda I_{2n}\right)=\det\left(\lambda^{2}I_{n}-A^{-1}C\right).$
Fix $\lambda>1$, it is enough to show that the $P(\lambda)\neq 0$, for which it is enough to prove that $\lambda^{2}I_{n}-A^{-1}C$ is positive definite. As $A\geq C$, we obtain that $I_{n}\geq A^{-1}C$, so $\lambda^{2}I_{n}-A^{-1}C=(\lambda^{2}-1)I_{n}+(I_{n}-A^{-1}C)$ is the sum of a positive definite and a positive semidefinite matrix, hence it is positive definite. ∎
Claim 4.4 and the definition of $\preceq$ immediately imply Corollary 1.13.
Corollary 1.13.
Let $(A_{i},B_{i},C_{i})\in\mathcal{G}(n)$, $i\in\{0,1\}$. Then
  1. (1)
    $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$Rightarrow $A_{1}-C_{1}\geq A_{0}-C_{0}$,
  2. (2)
    $A_{1}-C_{1}\geq A_{0}-C_{0}$ and $B_{1}-B_{0}$ is symmetric $\Rightarrow$ $(A_{0},B_{0},C_{0})\preceq(A_{1},B_{1},C_{1})$.
In order to prove Theorem 1.15 we need the following fact.
Fact 4.5.
If $(I_{n},B,I_{n})\in\mathcal{G}^{+}(n)$, then $B$ is symmetric.
Proof.
Assume that $(I_{n},B,I_{n})\in\mathcal{G}^{+}(n)$. By Claim 2.4 all positive eigenvalues of $M=iG\Omega$ are at most $1$. We can calculate the characteristic polynomial of $M$ as
$P(\lambda)=\det\left(M-\lambda I_{2n}\right)=\det\left[(\lambda^{2}-1)I_{n}-\lambda\frac{i}{2}\left(B-B^{\text{T}}\right)\right].$
It is clear that the constant term of $P$ is $P(0)=\det(-I_{n})=(-1)^{n}$, so by Vieta’s formula all roots of $P$ are ±$\pm 1$, both with multiplicity $n$ according to Theorem 2.3. This means that $P(\lambda)=(\lambda^{2}-1)^{n}$. Let $L$ be the coefficient of $\lambda^{2}$ in $P(\lambda)$ minus the coefficient of $\lambda^{2}$ in $(\lambda^{2}-1)^{n}$; clearly $L$ must be 00. Now we will calculate $L$ in another way. Let $N(\lambda)=(\lambda^{2}-1)I_{n}-\lambda(i/2)\left(B-B^{\text{T}}\right)$. In the determinant $P(\lambda)=\det(N(\lambda))$ if the term $\lambda^{2}$ comes only from the main diagonal, then it has the same coefficient as the coefficient of $\lambda^{2}$ in $(\lambda^{2}-1)^{n}$. As every non-diagonal element of $N(\lambda)$ is a constant multiple of $\lambda$, our only other option is to choose $-1$ from the main diagonal $n-2$ times, and if the remaining rows and columns are of number $j$ and $k$, then its contribution to $L$ is
$(-1)^{n-2}\left[-\frac{i}{2}(B_{jk}-B_{kj})\frac{i}{2}(B_{kj}-B_{jk})\right]=\frac{(-1)^{n-1}}{4}\left(B_{jk}-B_{kj}\right)^{2}.$
We need to sum these for all pairs $1\leq k<j\leq n$, which implies that
$L=\frac{(-1)^{n-1}}{4}\sum_{1\leq k<j\leq n}\left(B_{jk}-B_{kj}\right)^{2}.$
As $L=0$, we have $B_{jk}=B_{kj}$ for all $1\leq k<j\leq n$, so $B$ is symmetric. ∎
Now we can prove Theorem 1.15.
Theorem 1.15.
Let $(A_{i},B_{i},C_{i})\in\mathcal{G}(n)$, $i\in\{0,1\}$. The following are equivalent:
  1. (1)
    $(A_{0},B_{0},C_{0})\approx(A_{1},B_{1},C_{1})$;
  2. (2)
    $A_{1}-C_{1}=A_{0}-C_{0}$ and $B_{1}-B_{0}$ is symmetric.
Proof of Theorem 1.15.
[…]
$(A_{1}-A_{0}+rI_{n},B_{1}-B_{0},C_{1}-C_{0}+rI_{n})=(N+rI_{n},B_{1}-B_{0},N+rI_{n})\in\mathcal{G}^{+}(n).$
[…]
$(N+rI_{n},B_{1}-B_{0},N+rI_{n})\approx(I_{n},B_{1}-B_{0},I_{n}),$
so $(I_{n},B_{1}-B_{0},I_{n})\in\mathcal{G}^{+}(n)$. Thus Fact 4.5 implies that $B_{1}-B_{0}$ is symmetric. ∎
Acknowledgments
We are indebted to András Bodor, András Frigyik, Tamás Kiss, Mátyás Koniorczyk, Miklós Pintér, Géza Tóth and András Vukics for some illuminating conversations.
[…]